OpenStudy (anonymous):

write a parametrization for the equation given parametrically be the equations x=ln(2t) & y=2t^2 ?

OpenStudy (anonymous):

e^x=e^ln(2t)=>e^x=2t=>t=e^x/2 y=2(e^x/2)^2=e^2x/2